OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12

jigglypuff314 (jigglypuff314):

to do this, you'll want to get the absolute values by itself :) so divide both sides by Negative 3 what would you get?

OpenStudy (anonymous):

2x=6=4

OpenStudy (studygurl14):

−3|2x + 6| = −12 |2x + 6| = 4 2x + 6 = 4 Solve for x 2x + 6 = -(4) Solve for x

jigglypuff314 (jigglypuff314):

yep :) then |2x + 6| = 4 can be separated into 2x + 6 = 4 and 2x + 6 = -4 because when |a| = b a = b and a = -b

OpenStudy (anonymous):

x=5 x=-5

OpenStudy (anonymous):

????????!!!!!!!!!

OpenStudy (the_fizicx99):

−3|2x + 6| = −12 Divide both sides by negative 3 −3|2x + 6| = −12 /-3 /-3 |2x + 3| = 4 Ok so now you make a new equation and make the 4, -4 So, |2x + 3| = 4 & |2x + 3| = -4 Solve for x by subtracting both sides 3 Then dividing both sides by 2

jigglypuff314 (jigglypuff314):

close :) but not quite 2x + 6 = 4 subtract 6 from both sides -6 -6 2x = -2 then divide both sides by 2

OpenStudy (the_fizicx99):

Oh that typo, subtract 6 not 3 >.<

OpenStudy (anonymous):

x=-1 x=-5

jigglypuff314 (jigglypuff314):

correct! :D

OpenStudy (anonymous):

yea!!!!! thanx

jigglypuff314 (jigglypuff314):

glad we could help :)